How much must you deposit each year into your retirement account starting now and continuing through year 9 if you want to be able to withdraw $85,000 per year forever, beginning 26 years from now? Assume the account earns interest at 8% per year.

The amount to be deposited is determined to be $
Answer #1

R = 8% per year

If Funds of $85000 are withdrawn for perpetuity beginning 26 years from now,

Then,

Funds required at 26 years from now = 85000/8% = $1062500

Now, let, annual deposits for 9 years = P

Then,

1062500 = (P*(1.08^9 - 1)/.08)*1.08^16

P = (1062500/1.08^16)*(1/12.48755)

P = $24835.43

So, annual deposits for the first 9 years will be $24835.43 or $24835.